# AI-PACE: A Framework for Integrating AI into Medical Education

### v1

@@ Abstract
The integration of artificial intelligence (AI) into healthcare is accelerating, yet medical education has not kept pace with these technological advancements. This paper synthesizes current knowledge on AI in medical education through a comprehensive analysis of the literature, identifying key competencies, curricular approaches, and implementation strategies. The aim is highlighting the critical need for structured AI education across the medical learning continuum and offer a framework for curriculum development. The findings presented suggest that effective AI education requires longitudinal integration throughout medical training, interdisciplinary collaboration, and balanced attention to both technical fundamentals and clinical applications. This paper serves as a foundation for medical educators seeking to prepare future physicians for an AI-enhanced healthcare environment.
Keywords Medical Education $\cdot$ Artificial Intelligence $\cdot$ Curriculum Development $\cdot$ AI-PACE $\cdot$ Implementation Strategy
@@ 1 Introduction
Artificial intelligence’s (AI) role in medicine is rapidly changing; it is now an active clinical participant, from diagnostic imaging to administrative workflow optimization compared to an occasional aide in the past. This educational urgency is amplified by how rapidly consumers have adopted generative AI (GenAI). ChatGPT set the record for the fastest-growing user base, eclipsing previous record holders like TikTok and Instagram (Hu, 2023). With GenAI, patients are not waiting, they are forging ahead (Salmi, 2026; Stokel-Walker, 2024). They are already querying current models from OpenAI, Google, and Anthropic about medical conditions.
@@
OpenAI stated that in January of 2026, 230 million customers were asking ChatGPT about health and wellness per week (OpenAI, 2026). Seeing market demands, There are now two health focused GenAI tools for from the frontier labs, Claude for Healthcare (Anthropic, 2026) and ChatGPT Health (OpenAI, 2026). It is now possible for physicians to encounter patients who could activate AI voice assistants mid-visit to interpret recommendations in real-time. Hence, we are facing a "shadow" health system that clinicians must be prepared to navigate, whether or not formal training has equipped them to do so.
@@
Yet, current educational models fail to bridge the gap between the information age’s requirement for knowledge curation and the medical profession’s enduring need for deep biomedical competence (Wartman, 2019; Buja, 2019). While 44% of physicians report feeling inadequately prepared for new technologies (Minor, 2020), educational institutions have struggled to move beyond elective courses and short-term pilots. This disconnect presents a critical risk: a workforce expected to partner with AI tools without the foundational literacy to trust, verify, or oversee them.
Current educational efforts are often fragmented or siloed within specific domains. As evidenced by the short-term successes in specialty-specific pilots (van Kooten and others, 2024), such concentrated interventions risk rapid knowledge decay without longitudinal reinforcement. Similarly, Singla et al. (Singla and others, 2024) established a robust consensus on what competencies undergraduate medical students need but did not prescribe how these should evolve across the continuum of training. Other frameworks remain specialty-specific, such as Valikodath (Valikodath and others, 2021) defined for Ophthalmology.
This leaves a critical “Generalist Gap”: a lack of consensus on a comprehensive model for non-procedural or primary care physicians who will increasingly rely on AI-driven decision support. This paper addresses that structural gap. We move beyond the question of why AI should be taught to the practical challenge of how. By comparing existing models and synthesizing current best practices, we present the AI-PACE Framework (Psychomotor, Affective, Cognitive, Embedded).
@@
This model runs longitudinally, introducing AI literacy in the first year of medical school and continuing through practice. Consider that 79% of U.S. medical schools begin physical exam training within the first two months of the pre-clerkship phase to establish early competency (Uchida and others, 2019). The same logic applies here: early and repeated exposure to AI concepts builds the foundation physicians need to work effectively with these tools as they advance through training. Without this preparation, physicians risk becoming passive consumers of AI outputs rather than professionals equipped to evaluate and apply them appropriately.
@@ 2 Methods
To develop a comprehensive framework for AI in medical education, we conducted a systematic search and comparative analysis of the literature. We queried scholarly databases (PubMed, MEDLINE, ERIC) for peer-reviewed articles published between 2016 and 2025, using keywords related to artificial intelligence, medical education curricula, competency frameworks, and physician training. We also used Asta (Allen AI, 2025), from Allen AI to search for additional papers. This timeframe was selected to capture the rapid evolution of clinical AI applications and the corresponding educational responses.
Selection Criteria
Our initial search yielded 643 unique records. We screened these for relevance based on three inclusion criteria:
  1. 1.
    The proposal or evaluation of a structured AI curriculum or competency framework;
  2. 2.
    A target audience of medical students (Undergraduate Medical Education: UME), residents (Graduate Medical Education: GME), or practicing physicians (Continuing Medical Education: CME);
  3. 3.
    Publications in English.
We excluded purely technical papers, commentaries without curricular proposals, and studies focused on medical training outside of the United States. This process resulted in 23 key papers identified for in-depth review.
Analytical Approach
We performed a thematic analysis of the selected literature to identify recurring competencies and structural models. Frameworks were evaluated against three critical dimensions distilled from modern medical education theory:
  • Longitudinal Integration: Does the model support learning across the continuum of training (UME, GME, CME)?
  • Generalizability: Is the framework applicable to general medical practice, or is it restricted to specific specialties (e.g., Radiology)?
  • Domain Breadth: Does the framework integrate technical knowledge (Cognitive) with workflow application (Psychomotor) and professional attitudes (Affective)?
This comparative analysis (summarized in Table 1) revealed specific gaps in existing models, namely, the fragmentation of training into isolated workshops and the under-representation of the affective domain. The AI-PACE Framework operationalizes these competencies by grounding them in the Psychomotor, Affective, and Cognitive domains of Bloom’s Taxonomy (Anderson and Krathwohl, 2001), while adding “Embedded” as a structural pillar for longitudinal delivery.
@@
Study / Framework Target Audience & Scope Methodology Longitudinal Integration? Domains Covered Identified Gaps (vs. AI-PACE)
Valikodath et al. (Valikodath and others, 2021) (2021) Ophthalmology (UME to Fellowship) Perspective / Task Force Recommendations Yes (Proposed progression from med school to fellowship) Cog: Basics, Stats
Psycho: Surgical simulation
Aff: "Humanization" mentioned
Specialty-specific; lacks a generalized implementation model for all of medicine.
van Kooten et al. (van Kooten and others, 2024) (2024) Radiology (Residency) Pilot Study / Implementation No (3-day "fast-track" intensive course) Cog: Fundamentals
Psycho: Hands-on algorithm building
Ethics: Group discussion
Short-term intervention rather than longitudinal; specialty-specific; participants noted "hands-on" coding was less useful.
Singla et al. (Singla and others, 2024) (2024) UME (General Medical Students) Delphi Study (Expert Consensus) Suggested (Integration into existing curriculum recommended) Cog: Theory, Ethics, Law
Psycho: Application
Aff: Communication, Collaboration
Focused heavily on what to teach (competencies) rather than how to structure the educational continuum (UME->GME->CME).
Turner et al. (Turner et al., 2025) (2025) PCCM (Pulmonary & Critical Care) Perspective / Conceptual Partial (Mentions "Scalable Oversight") Cog: Ethics, Robustness
Psycho: Human-AI collaboration
Focuses on principles of safe AI integration rather than educational curriculum design; specialty-focused.
Masters (Masters, 2020) (2020) General (Admin & Educators) Conceptual Framework Conceptual (Curriculum mapping discussed) Cog: AI Literacy
Admin: Curriculum design
High-level conceptual guide for administrators; less focused on the practical syllabus for learners.
AI-PACE Framework General Medicine (UME, GME, CME) Integrative Review & Educational Theory Yes (Spiral curriculum across the full continuum) Psychomotor (Workflow)
Affective (Trust)
Cognitive (Knowledge)
Embedded (Longitudinal)
Addresses the Gap: Provides a unified, generalist model that spans the entire career continuum and explicitly operationalizes the Affective domain.
@@ 3 Current State of AI Education
Despite the proliferation of AI technologies in clinical practice, our analysis reveals a significant lag in the corresponding evolution of medical education frameworks. While recent efforts have begun to define necessary competencies, the landscape remains fragmented. A scoping review by Tolentino (Tolentino et al., 2024) found only two papers describing comprehensive curriculum frameworks for AI in medicine, with one focused specifically on ophthalmology. A recent scoping review by Rincón (Rincón and others, 2025) further confirmed these findings, revealing a lack of standardized AI curriculum frameworks and notable global discrepancies in implementation.
@@ 3.1 Fragmented Implementation and Knowledge Decay
Current AI education often relies on short-term, intensive "bootcamps" or elective workshops. For example, van Kooten (van Kooten and others, 2024) demonstrated the efficacy of a 3-day AI curriculum for radiology residents, which successfully improved confidence and self-reported knowledge in the short term. However, such concentrated burst of educational modules risk knowledge decay without reinforcement. They also can fail to integrate AI literacy into the daily clinical workflow where decision-making occurs. The majority of educational interventions consist of isolated courses, workshops, or lectures rather than integrated programs (Charow and others, 2021; Grunhut et al., 2021).
@@ 3.2 Specialty Imbalance and the Generalist Gap
Existing frameworks are heavily skewed toward image-centric specialties. Where AI education does exist, it focuses on primarily on radiology and imaging (Tolentino et al., 2024). We discussed previously the efforts focused on ophthalmology, and there have been critical principles outlined for Pulmonary and Critical Care Medicine (Turner et al., 2025). While specialty-specific models offer excellent depth, AI-PACE provides the necessary transferability for general medical education. By focusing on universal competencies, such as critical appraisal of algorithms and appropriate trust calibration, the framework prepares generalist physicians to evaluate AI-driven decision support tools across any clinical domain, rather than focusing on image-specific interpretation.
3.3 The "Human" Gap: Affective Competencies
Most critically, current models frequently overlook the "human" element of human-AI interaction. Using a Delphi approach provided a strong consensus on what to teach (Singla and others, 2024), validating competencies in ethics, theory, and application, but that particular framework focuses primarily on knowledge and skills. What remains under-addressed is the Affective Domain: the calibration of trust, the recognition of automation bias, and the preservation of empathy in an AI-mediated patient encounter. Rincón highlighted an emerging educational need for strengthening transversal skills, such as ethical awareness, to support the integration of AI as a practical tool rather than a standalone subject (Rincón and others, 2025).
@@ 4 The AI-PACE Framework
Figure 1: AI-PACE Taxonomic Framework for Medical AI Competencies within Bloom’s Taxonomy Domains.
@@ 4.2 Affective Domain (Appropriate Entrustment & Trust Calibration)
The affective domain encompasses the attitudes, values, and emotional intelligence required for human-AI partnership. This domain is critical for addressing the "human gap" identified in our analysis. However, framing this as "trust" undersells what clinical training actually requires. Clinicians must learn to assess AI outputs as they would a junior colleague’s work. This means understanding where the tools can fail (i.e. LMMs summarize effectively but struggle to extract qualitative meaning from patient narratives). Therefore, curricula must address these boundaries explicitly rather than treating AI competence as a binary trust decision.
  • Appropriate Trust Calibration: Developing a balanced perspective on AI capabilities to avoid both "automation bias" (over-reliance) and "algorithm aversion" (unwarranted skepticism) (Grunhut et al., 2022).
@@ 4.3 Cognitive Domain (Knowledge)
  • Fundamentals of AI: Understanding basic concepts, terminology (machine learning, deep learning), and knowing that ML models generate probabilistic predictions based on patterns, not rigid rules (Masters, 2020; Paranjape et al., 2019). Foundational training must also update outdated skepticism. The 2022-era critique of AI as "stochastic parrots", probabilistic word repeaters, no longer reflects current model capabilities (OpenAI, 2025; Google, 2025; Anthropic, 2025). Reasoning and agentic functions have advanced substantially. Curricula should ground skepticism in actual technical limitations.
@@ 4.4 Embedded (Longitudinal Integration)
The "Embedded" pillar addresses the structural limitations of fragmented "bootcamp" models by distributing AI competencies across the full medical training continuum, which aligns with the higher levels of Harden’s “integration ladder” (Harden, 2000). Rather than treating AI as an isolated elective or a standalone "AI week," this approach aligns with Han’s work (Han and others, 2019) by weaving AI literacy into the existing fabric of pre-clerkship and clerkship coursework.
This longitudinal model directly bridges the "Generalist Gap" by establishing foundational cognitive and affective competencies early in training, which then branch into specialized psychomotor skills during residency and independent practice.
Implementation must also address the faculty gap. Early-career faculty often have strong interest in AI literacy but may hesitate to enter classes where students could outpace them. This creates a disconnect: the faculty best positioned to integrate emerging tools might also be those most wary of the optics of learning along with their students. Therefore, parallel upskilling pathways for junior faculty are essential, without them, we cannot expect faculty to model the critical evaluation and entrustment behaviors the curriculum intends to teach.
@@
  • Graduate Medical Education (GME): Shifts toward specialty-specific applications and workflow integration. Residents move from theory to clinical utility, practicing "human-in-the-loop" verification during specialty rotations (e.g., Internal Medicine or Radiology clerkships) through hands-on, case-based learning.
@@
To operationalize this approach, Table 2 provides a curriculum map that aligns the AI-PACE domains with specific milestones and existing integration points across the medical education continuum. This map transitions from foundational theory in undergraduate years to advanced clinical oversight and leadership in continuing practice.
@@
Training Phase Primary Focus Cognitive (Knowledge) Psychomotor (Skills/Workflow) Affective (Attitudes/Trust) Integration Point & Learning Approach
Undergraduate (UME - Pre-clinical) Foundations AI Terminology & Concepts; ML vs. Deterministic logic; Probabilistic nature of AI; Data ethics & bias. Basic prompt engineering for study aids; Navigating literature search AI; Overview of AI applications. Recognizing automation bias; Valuing data privacy; Skepticism of "Black Box." Introductory courses & integrated modules in Epidemiology, Biostatistics, & Health Systems Science.
Undergraduate (UME - Clinical) Exposure Interpreting AI metrics (AUC, Sensitivity); Regulatory landscapes (FDA); Data Science Basics. Critical appraisal of AI tools during rounds; Using AI for differential diagnosis; Communicating results to patients. Calibration of trust; Empathy in AI-mediated patient encounters. Clinical Clerkships (e.g., IM, Radiology) via integrated modules and clinical exposure.
Graduate (GME - Residency) Application Specialty-Specific AI Applications; Sepsis predictors; Image triage algorithms. Clinical Integration & Decision Making; "Human-in-the-loop" verification; EHR documentation; Error reporting. Professional accountability; Collaboration with data science teams. Hands-on applications & case-based learning during Morning Reports, Grand Rounds, & QI projects.
Continuing (CME - Practice) Mastery & Leadership Emerging LLM capabilities; Quality & Performance Assessment; Liability updates. Workflow Optimization; Selecting/implementing AI tools; Leadership in AI Implementation. Fostering a culture of safety; Mentoring junior staff on AI oversight. Targeted workshops, just-in-time training, and institutional leadership roles.
This outlines the Embedded (Longitudinal) pillar of AI-PACE by mapping the Cognitive, Psychomotor, and Affective domains to specific milestones in training.
Integrated Pedagogical Methods:
  • Case-Based Learning: Use specialty-specific "AI-Fail" case studies to teach trust calibration and verification.
  • Simulation Objective Structured Clinical Examination (OSCEs): Practice explaining AI-generated diagnoses and algorithmic limitations to standardized patients.

### v2

@@ Abstract
Medical AI education remains fragmented, specialty-skewed, and lacks longitudinal structure, particularly for generalist physicians. Through an integrative review of 23 peer-reviewed articles (2016–2025), we identified three structural gaps: short-term interventions without reinforcement, procedural-field bias, and consistent under-representation of the Affective domain. We present AI-PACE (Psychomotor, Affective, Cognitive, Embedded), a Bloom’s Taxonomy-grounded framework organizing AI competencies longitudinally across undergraduate, graduate, and continuing medical education.
Keywords Medical Education $\cdot$ Artificial Intelligence $\cdot$ Curriculum Development $\cdot$ Implementation Strategy
@@ 1 Introduction
Artificial intelligence’s (AI) role in medicine is rapidly changing; it is now an active clinical participant, from diagnostic imaging to administrative workflow optimization, compared to an occasional aide in the past. This educational urgency is amplified by how rapidly consumers have adopted generative AI (GenAI). When released in November of 2022, ChatGPT set the record for the fastest-growing user base, eclipsing previous record holders like TikTok and Instagram (Hu, 2023). With GenAI, patients are not waiting for clinicians, they are forging ahead of healthcare adoption (Salmi, 2026; Stokel-Walker, 2024). They are already querying current models from OpenAI, Google, and Anthropic about medical conditions.
@@
OpenAI stated that in January of 2026, 230 million customers were asking ChatGPT about health and wellness per week (Silberling, 2026). Both OpenAI and Anthropic have now released health-focused products in reaction to these market demands (Claude for Healthcare (Anthropic, 2026) and ChatGPT Health (OpenAI, 2026)). It is now possible for physicians to encounter patients who could activate AI voice assistants mid-visit to interpret recommendations in real-time. Hence, we are facing a “shadow” health system that clinicians must be prepared to navigate, whether or not formal training has equipped them to do so.
@@
Yet, current educational models fail to bridge the gap between the information age’s requirement for knowledge curation and the medical profession’s enduring need for deep biomedical competence (Wartman, 2019; Buja, 2019). While 44% of physicians report feeling inadequately prepared for new technologies (Minor, 2020), educational institutions have struggled to move beyond elective courses and short-term pilots. Existing efforts tend to be fragmented, specialty-specific pilots which show short-term gains but risk knowledge decay. Without longitudinal reinforcement (van Kooten and others, 2024), competency frameworks identify what to teach but not how skills should evolve across training (Singla and others, 2024), and most training approaches remain anchored to procedural specialties like radiology or ophthalmology (Valikodath and others, 2021). The result is a workforce expected to partner with AI tools without the foundational literacy to evaluate, verify, or oversee them.
This leaves a “Generalist Gap”: a lack of consensus on a comprehensive model for non-procedural or primary care physicians who will increasingly rely on AI-driven decision support. This paper addresses that structural gap. We move beyond the question of why AI should be taught to the practical challenge of how. To that end, we first clarify several terms that are often conflated in the literature. Artificial intelligence (AI) refers broadly to computational systems designed to perform tasks that typically require human cognition. Machine learning (ML) is a subset of AI in which models learn patterns from data rather than following explicitly programmed rules. Generative AI (GenAI) describes a class of ML models, including large language models (LLMs), capable of producing novel text, images, or other outputs. Large language models (LLMs) are a specific GenAI architecture trained on large text corpora to generate and interpret natural language. These distinctions matter for curriculum design: the competencies required to use a diagnostic imaging algorithm differ meaningfully from those needed to evaluate an LLM-generated clinical summary.
@@
By comparing existing models and synthesizing current best practices, we present the AI-PACE Framework (Psychomotor, Affective, Cognitive, Embedded). This model runs longitudinally, introducing AI literacy in the first year of medical school and continuing through practice. Consider that 79% of U.S. medical schools begin physical exam training within the first two months of the pre-clerkship phase to establish early competency (Uchida and others, 2019). The same logic applies here: early and repeated exposure to AI concepts builds the foundation physicians need to work effectively with these tools as they advance through training. Without this preparation, physicians risk becoming passive consumers of AI outputs rather than professionals equipped to evaluate and apply them appropriately.
@@ 2 Methods
To develop a comprehensive framework for AI in medical education, we conducted an integrative review of the literature. Integrative reviews, as defined by Whittemore and Knafl (Whittemore and Knafl, 2005), are specifically designed to synthesize diverse methodological approaches, including empirical and conceptual work, to build new theoretical frameworks. This methodology was selected given our primary objective: not to evaluate the effectiveness of existing curricula, but to map out existing AI competency frameworks and identify structural gaps that a new model could address.
2.1 Search Strategy
We searched PubMed, MEDLINE, and ERIC using the following query: (“artificial intelligence,” “machine learning,” “AI,” “data science,” “deep learning”), medical education terms (“medical education,” “medical student*,” “residency,” “physician education,” “continuing medical education”), and curriculum/competency terms (“curriculum,” “curricula,” “competenc*,” “training program,” “educational program”).
The search was limited to publications from 2016 to 2025 to capture the rapid evolution of AI applications in healthcare and the corresponding educational responses. We also queried Asta (Allen AI, 2025) to identify papers our initial query might have missed.
Database selection prioritized clinical pedagogy and health professions education literature. Scopus, Web of Science, and IEEE Xplore were not included because our focus was on clinical and educational implementation rather than technical AI development.
2.2 Study Selection
The initial search yielded 643 records. We applied six inclusion criteria: (1) focus on AI curriculum or competency frameworks, (2) medical education setting, (3) publication between 2016–2025, (4) English language, (5) peer-reviewed, and (6) focus on US medical programs. The decision to focus on US programs was pragmatic: AI-PACE was developed with the explicit intent of building a pilot course for medical students in the United States, and mapping competencies to the structural phases of US training (UME, GME, CME) was essential for that application. Exclusion criteria included conference abstracts without full text, studies focused on patient care rather than education, and purely technical AI papers without an educational component. Article screening was conducted by a single reviewer (S.P.M.). This is a limitation of the current work and is addressed in the Limitations section. Following full-text review, 23 articles met all inclusion criteria.
2.3 Data Analysis
We performed thematic analysis of the 23 papers to identify recurring competencies and curriculum structures. Documents were analyzed for themes across AI competencies, curricular frameworks, educational approaches, and implementation challenges. Competencies were organized across the three domains of Bloom’s Taxonomy following Charow et al. (Charow and others, 2021). Curriculum development stages were mapped to established medical education frameworks (Harden, 1986; Thomas et al., 2022).
This analysis (summarized in Table 1) revealed specific gaps in existing models, namely, the fragmentation of training into isolated workshops and the under-representation of the affective domain. The AI-PACE Framework operationalizes these competencies by grounding them in the Psychomotor, Affective, and Cognitive domains of Bloom’s Taxonomy (Anderson and Krathwohl, 2001), while adding “Embedded” as a structural pillar for longitudinal delivery.
@@ 2.3 Data Analysis
Study / Framework Audience & Scope Methodology Longitudinal Integration? Domains Covered Gaps Identified
Valikodath et al., 2021 (Valikodath and others, 2021) Ophthalmology (UME to Fellowship) Perspective / Task Force Recommendations Yes (Proposed progression from med school to fellowship) Fundamentals, surgical simulation, humanization Specialty-specific; no generalist model
van Kooten et al., 2024 (van Kooten and others, 2024) Radiology (Residency) Pilot Study / Implementation No (3-day “fast-track” intensive course) Fundamentals, algorithm building, group ethics Short-term, specialty-specific; no longitudinal reinforcement
Singla et al., 2024 (Singla and others, 2024) UME (General Medical Students) Delphi Study (Expert Consensus) Suggested (Integration into existing curriculum recommended) Theory, ethics, law, application, communication Defines what to teach; does not address continuum progression
Turner et al., 2025 (Turner et al., 2025) PCCM (Pulmonary & Critical Care) Perspective / Conceptual Partial (Mentions “Scalable Oversight”) Ethics, robustness, human-AI collaboration Safe integration principles; not a curriculum design framework
Masters, 2020 (Masters, 2020) General (Admin & Educators) Conceptual Framework Conceptual (Curriculum mapping discussed) AI literacy, curriculum design Administrator-focused; limited learner-facing guidance
AI-PACE Framework General Medicine (UME, GME, CME) Integrative Review & Bloom’s Taxonomy Yes (Spiral curriculum across the full continuum) Psychomotor (Workflow); Affective (Trust); Cognitive (Knowledge); Embedded (Longitudinal) Addresses the gap: Generalist, continuum-spanning; explicitly includes Affective domain
@@ 3 Current State of AI Education
Despite the proliferation of AI technologies in clinical practice, our analysis reveals a significant lag in the corresponding evolution of medical education frameworks. While recent efforts have begun to define necessary competencies, the field remains fragmented. A scoping review by Tolentino (Tolentino et al., 2024) found only two papers describing comprehensive curriculum frameworks for AI in medicine, with one focused specifically on ophthalmology. A recent scoping review by Rincón (Rincón and others, 2025) further confirmed these findings, revealing a lack of standardized AI curriculum frameworks and notable global discrepancies in implementation.
@@ 3.1 Fragmented Implementation and Knowledge Decay
Current AI education often relies on short-term, intensive “bootcamps” or elective workshops. For example, van Kooten (van Kooten and others, 2024) demonstrated the efficacy of a 3-day AI curriculum for radiology residents, which successfully improved confidence and self-reported knowledge in the short term. However, such concentrated bursts of educational modules risk knowledge decay without reinforcement. They also can fail to integrate AI literacy into the daily clinical workflow where decision-making occurs. The majority of educational interventions consist of isolated courses, workshops, or lectures rather than integrated programs (Charow and others, 2021; Grunhut et al., 2021).
Recent interventions have begun to demonstrate measurable gains. Levingston et al. (Levingston et al., 2024) developed an AI literacy course for first-year medical students integrating foundational concepts, practical tool use, and ethics. Hopson et al. (Hopson and others, 2025) evaluated a structured 4-week pre-medical AI curriculum using a pre-post design and found significant improvements in AI knowledge scores. These results are encouraging, but both interventions are single-site and lack longitudinal follow-up, reinforcing the need for integrated, continuum-spanning approaches.
@@ 3.2 Specialty Imbalance and the Generalist Gap
Existing frameworks are heavily skewed toward image-centric specialties. Where AI education does exist, it focuses primarily on these fields (Tolentino et al., 2024). We discussed previously the efforts focused on ophthalmology, and there have been important principles outlined for Pulmonary and Critical Care Medicine (Turner et al., 2025). Specialty-specific models address depth within defined domains. AI-PACE is designed to address the transferability gap, preparing generalist physicians to evaluate AI-driven decision support across specialties rather than within a single domain.
3.3 The “Human” Gap: Affective Competencies
Current models also tend to under-address the human element of human-AI interaction. Using a Delphi approach provided a strong consensus on what to teach (Singla and others, 2024), validating competencies in ethics, theory, and application, but that particular framework focuses primarily on knowledge and skills. What remains under-addressed is the Affective Domain: the calibration of trust, the recognition of automation bias, and the preservation of empathy in an AI-mediated patient encounter. Rincón highlighted an emerging educational need for strengthening transversal skills, such as ethical awareness, to support the integration of AI as a practical tool rather than a standalone subject (Rincón and others, 2025).
@@ 4 The AI-PACE Framework
Figure 1: The AI-PACE Taxonomic Framework organizes medical AI competencies across three domains of Bloom’s Taxonomy: Cognitive (knowledge), Psychomotor (skills), and Affective (attitudes); with specific competencies mapped to each domain. Adapted from Charow et al. (Charow and others, 2021), Grunhut et al. (Grunhut et al., 2021), Masters (Masters, 2020), and Li et al. (Li et al., 2019).
@@ 4.2 Affective Domain (Appropriate Entrustment & Trust Calibration)
The affective domain encompasses the attitudes, values, and emotional intelligence required for human-AI partnership. This domain is central to addressing the “human gap” identified in our analysis. However, framing this as “trust” undersells what clinical training actually requires. Clinicians must learn to assess AI outputs as they would a junior colleague’s work. This means understanding where the tools can fail (i.e., LLMs summarize effectively but struggle to extract qualitative meaning from patient narratives). Therefore, curricula must address these boundaries explicitly rather than treating AI competence as a binary trust decision.
Trust calibration in this context carries a duality that curricula must acknowledge. Clinicians must learn to evaluate the trustworthiness of AI outputs (trust in the AI), but educators must simultaneously assess whether the trainee has developed sufficient clinical judgment to make that evaluation reliably (trust in a trainee’s competencies). Gin et al. propose applying the Entrustable Professional Activity (EPA) framework to AI tools as trustees, assessing AI across ability, integrity, and benevolence in parallel with how we assess trainees, as a structured approach to operationalizing both sides of this trust relationship (Gin and others, 2025).
  • Appropriate Trust Calibration: Developing a balanced perspective on AI capabilities to avoid both “automation bias” (over-reliance) and “algorithm aversion” (unwarranted skepticism) (Grunhut et al., 2022).
@@ 4.3 Cognitive Domain (Knowledge)
  • Fundamentals of AI: Understanding basic concepts, terminology (machine learning, deep learning), and knowing that ML models generate probabilistic predictions based on patterns, not rigid rules (Masters, 2020; Paranjape et al., 2019). Foundational training must also update outdated skepticism. The 2022-era critique of AI as “stochastic parrots”—probabilistic word repeaters—no longer reflects current model capabilities (Qiu and others, 2025; McCoy and others, 2025; Anthropic, 2025; Google DeepMind, 2025). Reasoning and agentic functions have advanced substantially. Curricula should ground skepticism in actual technical limitations.
@@ 4.4 Embedded (Longitudinal Integration)
The “Embedded” pillar addresses the structural limitations of fragmented bootcamp models by distributing AI competencies across the full medical training continuum, which aligns with the higher levels of Harden’s “integration ladder” (Harden, 2000). Rather than treating AI as an isolated elective or a standalone “AI week,” this approach aligns with Han’s work (Han and others, 2019) by weaving AI literacy into the existing fabric of pre-clerkship and clerkship coursework. This longitudinal model directly bridges the “Generalist Gap” by establishing foundational cognitive and affective competencies early in training, which then branch into specialized psychomotor skills during residency and independent practice.
Implementation must also address the faculty gap. Early-career faculty often have strong interest in AI literacy but may hesitate to enter classes where students could outpace them. This creates a disconnect: the faculty best positioned to integrate emerging tools might also be those most wary of the optics of learning along with their students. Therefore, parallel upskilling pathways for junior faculty are essential; without them, we cannot expect faculty to model the critical evaluation and entrustment behaviors the curriculum intends to teach.
@@
  • Graduate Medical Education (GME): Shifts toward specialty-specific applications and workflow integration. Residents move from theory to clinical utility, practicing “human-in-the-loop” verification during specialty rotations (e.g., Internal Medicine or Radiology clerkships) through hands-on, case-based learning.
@@
To operationalize this approach, Table 2 provides a curriculum map that aligns the AI-PACE domains with specific milestones and existing integration points across the medical education continuum. This map transitions from foundational theory in undergraduate years to advanced clinical oversight and leadership in continuing practice. It outlines the Embedded (Longitudinal) pillar of AI-PACE by mapping the Cognitive, Psychomotor, and Affective domains to specific milestones in training.
@@
Training Phase Primary Focus Cognitive (Knowledge) Psychomotor (Skills/Workflow) Affective (Attitudes/Trust); Integration Point & Learning Approach
Undergraduate Foundation (UME — Pre-Clinical) Foundations AI terminology & concepts; ML vs. deterministic logic; probabilistic nature of AI Basic prompt engineering for study aids; navigating literature search Recognizing automation bias; valuing data privacy; skepticism of “black boxes.”
Introductory courses & integrated modules in Epidemiology, Biostatistics, & Health Systems Science.
Undergraduate (UME — Clinical) Exposure Interpreting AI metrics (AUC, sensitivity); regulatory landscapes (FDA); data science basics Critical appraisal of AI tools during rounds; using AI for differential diagnosis; communicating results to patients Calibration of trust; empathy in AI-mediated patient encounters.
Clinical clerkships (e.g., IM, Radiology) via integrated modules and clinical exposure.
Graduate (GME — Residency) Application Specialty-specific AI applications; sepsis predictors; image triage algorithms Clinical integration & decision making; “human-in-the-loop” verification; EHR documentation; error reporting Professional accountability; collaboration with data science teams.
Hands-on applications & case-based learning during morning reports, grand rounds, & QI projects.
Continuing (CME — Practice) Mastery & Leadership Emerging LLM capabilities; quality & performance assessment; liability updates Workflow optimization; selecting/implementing AI tools; leadership in AI implementation Fostering a culture of safety; mentoring junior staff on AI oversight.
Targeted workshops, just-in-time training, and institutional leadership roles.
4.5 Integrated Pedagogical Methods
  • Case-Based Learning: Use specialty-specific “AI-Fail” case studies to teach trust calibration and verification.
  • Simulation Objective Structured Clinical Examinations (OSCEs): Practice explaining AI-generated diagnoses and algorithmic limitations to standardized patients.